Show that cos³θ +sin³θ =1/4(cos³θ + 3cosθ - sin³θ + 3sinθ)



1
Expert's answer
2021-05-07T09:26:18-0400
  • It is a rearrangement of several trigonometric identities.
  • As it is known,

"\qquad\qquad
\begin{aligned}
\small \sin3\theta &=\small 3\sin\theta-4\sin\theta^3\\
\small \cos3\theta&=\small 4\cos^3\theta -3\cos\theta
\end{aligned}"

  • Then by rearranging,

"\qquad\qquad
\begin{aligned}
\small 4\cos^3\theta &=\small \cos3\theta+3\cos\theta\\
\small \cos^3\theta&=\small \frac{1}{4}\cdot(\cos3\theta+3\cos\theta)\cdots\cdots(1)\\
\\
\small 4\sin^3\theta&=\small 3\sin\theta-\sin^3\theta\\
\small \sin^3\theta&=\small \frac{1}{4}\cdot(3\sin\theta-\sin^3\theta)\cdots\cdots(2)\\
\\
&\small\text{Then by (1)+(2)}\\
\\
\small \sin^3\theta+\cos^3\theta &=\small \frac{1}{4}(3\sin\theta-\sin^3\theta) +\frac{1}
{4}(\cos3\theta+3\cos\theta)\\
&=\small \frac{1}{4}\Big[3\cos\theta+\cos3\theta+3\sin\theta-\sin^3\theta\Big]
\end{aligned}"